OpenStudy (anonymous):

An American Red Cross plane needs to drop emergency supplies to victims of a hurricane in a remote part of the world. The plane is traveling horizontally at 100.0 m/s at a height of 50.0 m above the ground. What horizontal distance does the package travel before striking the ground?

OpenStudy (anonymous):

Find the time needed by the package to strike the ground first \(t=\sqrt{\frac{2h}{g}}\) then find horizontal distance with \(x=v_xt\)

OpenStudy (anonymous):

So would the t=3? then giving you x=(100m/s)(3s). x=300m. Right?

OpenStudy (anonymous):

Corrrrrect :D

OpenStudy (anonymous):

Thanks :)